• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jwexler

    Log entry time 00:01:35 on June13,2010

    Entry number 326122

    keyword=Shift Summary 20:00-00:00

    Shift Summary: 8PM-12AM

    20:00 - Carbon production continues.
    20:30 - Trouble controlling HVs for the right arm. Telnet shows 'packed' mode on the HV controller, HV gui not working appropriately (?)
    21:00 - HV GUI works fine, just takes some time to start up.
    21:05 - North LINAC went down.
    22:55 - North LINAC back up, currently starting beam recovery on PbD #1.
    23:00 - MCC says Hall B isn't seeing any beam. Taking beam away from us to sort that out.
    23:22 - MCC ready to send CW beam again. Restarting beam recovery.
    00:00 - Beam recovery in progress. Putting in PbD#1, proceeding to position scan for frame.